How do you solve this linear system using substitution?
3/8x + 3/4y = 12
2/3x + 1/2y = 13

Answer by Alan3354(69443)

How do you solve this linear system using substitution?
3/8x + 3/4y = 12
2/3x + 1/2y = 13
----------------
Step 1, get rid of fractions
3x + 6y = 96 eqn 1 times 8
x + 2y = 32 then divided by 3
4x + 3y = 78 eqn 2 times 6
--------------
x + 2y = 32 Solve for x
x = 32 - 2y
--------------
Sub for x into the other eqn
4(32-2y) + 3y = 78
128 - 8y + 3y = 78
-5y = -40
y = 8
------------
x = 32 - 2y
x = 32 - 2*8
x = 16
That's how.